OpenStudy (anonymous):

Help with a few algebra questions? (:

OpenStudy (anonymous):

\[f(x)=7\times0.32^{x}\]

OpenStudy (australopithecus):

go on?

OpenStudy (anonymous):

\[g(x)=1.3\left( \frac{ 1 }{ 4 } \right)^{x}\]

OpenStudy (australopithecus):

what is the question?

OpenStudy (anonymous):

Oh tell whether the function represents exponential growth or exponential decay, Identify the growth or decay factor

OpenStudy (anonymous):

Where a = initial amount b = growth/decay factor x = time y = ending amount Growth When a > 0 and b > 1, the function models growth. (b is called the growth factor) (a represents the initial amount) Decay When a > 0 and 0 < b < 1, the function models decay. (b is called the decay factor) (a represents the initial amount)

OpenStudy (anonymous):

I'm so confused

OpenStudy (australopithecus):

When the exponent is negative that is decay https://www.wolframalpha.com/input/?i=y+%3D+10^-x When the exponent is positive that is growth https://www.wolframalpha.com/input/?i=y+%3D10^x

OpenStudy (australopithecus):

make sense?

OpenStudy (anonymous):

Not really. :/

OpenStudy (anonymous):

The formula for exponential growth is \[y=a*b^x\]So in your first question you have \[f(x)=7*.32^x\]so your a = 7 your b = .32 Now for exponential GROWTH your "a" value must be greater than 0 so a > 0 and for your "b" value for exponential growth it must be greater than 1 so b > 1 But for exponential DECAY a > 0 and b is between 0 and 1 So in your first question you have a= 7 which is greater than 0 but your b value is 0.32 which is between 0 and 1 therefore you have exponential DECAY Try your other question using this same principle and tell me what you think it is Hope that helped